2A bird of mass 0.6 kg stands on the midpoint of a horizontal rope. The rope sags and
angle of 150o is formed.
Find the tension in the rope.
Ans: T = 11.6 N
1
Expert's answer
2020-05-14T09:01:59-0400
"mg=2T\cos(\theta/2)"

"T=\frac{mg}{2\cos(\theta/2)}=\frac{0.6\times 9.8}{2\cos(150^{\circ}/2)}=11.36\:\rm N"